# Leveraging Discrete Function Decomposability for Scientific Design

### v1

@@ 2.1 Formal exposition of DADO
The goal of DADO is to obtain a generative model, $p_{\theta}(x)$, that maximizes the EDA objective, $\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]$, while leveraging decomposability in $f(x)$ for optimization efficiency. DADO requires as input a decomposed version of $f(x)$, which can be described by an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. As noted in the introduction and shown in our experiments on proteins, identifying useful decomposability is feasible in practice. Given the junction tree decomposition, we write $f(x)=\sum_{i\in\mathcal{N}}f_{i}(\tilde{x}_{i})+\sum_{(i,j)\in\mathcal{E}}f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$, where we refer to $f_{i}(\tilde{x}_{i})$ as “node component functions” and $f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$ as “edge component functions” (Fig. 1a). This functional structure is intimately related to notions of the “epistatic landscape” of a protein’s property function and spectral analysis (see Sec. A.1.1). When the component functions are not known a priori, they can be parameterized and fit to labeled data.
@@ 2.1.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models (pearl88_bp_textbook; shah_14_mit_inference-algorithms_message-passing), it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ (vlassis04_maxplus). In particular, message-passing can be used to find a global maximum of a function defined on an undirected junction tree, $\mathcal{T}$, by making use of its topology for optimal time-complexity. In particular, one takes the junction tree, which is undirected, and roots it to obtain a directed tree, which induces a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
@@ 2.1.2 From classical message passing to distributional optimization
In the same way that an EDA transforms $\arg\max_{x}f(x)$ into a distributional optimization problem, we can rewrite the equivalent message-passing objective in Eq. 5 as DO Because the original optimization problems over $x$ are equivalent, their DO formulations are equivalent too,
$\arg\max\nolimits_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]=\arg\max\nolimits_{\theta}\Big(\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\Big),$ (6)
@@
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. We emphasize that DADO’s update is more statistically efficient than a naive EDA’s because DADO gets to use all $K$ samples for weighted MLE on each lower-dimensional factor distribution (i.e., ratio of number of samples to number of dimensions is larger). Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (steps 1 and 2 illustrated in Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics (brookes20_eda_em), so we include it in Alg. 1 on lines 16 and 18. Choosing $W$ to be the identity recovers our derivation above. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 10) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.4).
@@
1:Given: Decomposition $\left(\text{junction tree}\;\mathcal{T}=(\mathcal{N},\mathcal{E}),\;\text{component functions}\;\{f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})\}_{(i,j)\in\mathcal{E}}\;\text{and}\;\{f_{i}(\tilde{x}_{i})\}_{i\in\mathcal{N}}\right)$
@@
13:  end for
14:  for all $i\in\mathcal{N}$, in parallel do $\triangleright$ Update each search distribution factor using value functions as weights
@@ 3 Related work
The closest work to ours in the EDA community is the Factorized Distribution Algorithm (FDA), wherein a decomposition is leveraged to factorize the conditional probability table (CPT) that parameterizes the search distribution (muhlenbein99_fda). However, FDA cannot leverage message-passing to give different weights to different variables so as to more efficiently update the search distribution. In related work, pelikan05_boa replaced the CPTs with Bayesian networks, but still had no means to coordinate among variables through message-passing or the like.
In a complementary line of work on policy optimization in reinforcement learning (analogous to learning the search distribution in EDAs), coordination between variables is enabled by message-passing on a factorized search distribution/policy, but this line of work is only suitable for graph topologies that are Markov chains—that is, chain graphs. These graphs arise from reward functions, $f(x)$, that decompose in time according to a first-order Markov assumption (peters07_rwrpeng19_awrnair20_awac). DADO can be viewed as generalizing policy optimization over chains to arbitrary graphs. We explore this connection in more detail in Sec. A.4.
Adjacent to our problem of interest, because they focus solely on real-valued design spaces,  grudzien24_fgm introduces a new way to discover a function decomposition from data and then demonstrates that using it improves out-of-distribution generalization and thus, downstream optimization performance. Separately, the Bayesian optimization (BO) community has developed methods for dynamically inferring a function decomposition from data during active learning. Although they make use of classical message-passing for coordination, they do not employ distributional optimization, nor can these methods be readily generalized to do so (kandasamy15_additive-BOhan21_tree-BOhoang18_JT-BOrolland18_JT-BObardou24_loopy_BO). This community has shown that using approximate or even random decompositions can be helpful (ziomek23_random-decompositions_BO). DADO could, in principle, be used within the BO inner loop, although such an investigation is beyond the scope of the present work.
@@ 4 Experimental results
We perform two sets of experiments with increasing resemblance to real-world scientific design. In each setting we compare a naive EDA, which is unaware of the function decomposition, to DADO, which is aware of it. In the first setting, we create synthetic functions $f(x)$, while in the second setting, we focus on functions derived from real protein data. In all experiments we designed sequences of fixed length $L$, where each position is one of $D=20$ amino acids. We used $N=100$ EDA training iterations and $G=1$ gradient steps for each weighted maximum likelihood update. For the search distribution of DADO, we used MLP-based autoregressive neural networks for each node—specifically, each variable $x_{j}$ in $\tilde{x}_{i}$ is modeled by an MLP of size $[64,64]$ that takes as input all conditioning variables (i.e., $\tilde{x}_{\text{p}(i)}$ and any preceding variables in $\tilde{x}_{i}$) and outputs $D$ logits. Within $\tilde{x}_{i}$, an arbitrary decoding order was used, and any overlapping variables, that is, $x_{j}$ in both $\tilde{x}_{i}$ and $\tilde{x}_{\text{p}(i)}$, were considered the purview of the parent and thus omitted from the child except as conditioning. Our naive EDA implementation is the same, only assuming all variables are in one meta-variable—hence, a standard autoregressive model over all variables. For gradient descent to fit the search distribution, we used the AdamW optimizer (loshchilov17_adamw) with default momentum parameters $\beta_{1}=0.9,\beta_{2}=0.999$. All code is implemented using Jax, Flax, and Optax (deepmind20_jax). For the shaping function we choose ${W}(s)=\exp\frac{s}{\beta}$ and tuned the temperature, $\beta$. To choose $\beta$ and the learning rate, $\eta$, we performed a hyperparameter sweep separately for each method and each $f(x)$, taking $(\eta^{*},\beta^{*})$ with the largest sample mean at the last iteration. For synthetic experiments, we swept 100 combinations between $\eta\in[10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.1,\;8]$. For protein experiments, we swept only 54 pairs between $\eta\in[5\times 10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.01,\;5]$, for expediency. We confirmed that neither method’s $\eta^{*}$ or $\beta^{*}$ was on a boundary of the sweep. $(\eta^{*},\beta^{*})$ were then used for all replicate runs over random seeds dictating the initial search distribution and sampling. Although not essential to our method as described in Alg. 1, in our implementation we reduce the variance of the search distribution update by using mean-shifted $Q$-functions, $Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})-\mathbb{E}_{\{\tilde{x}_{i}^{k}\}}[Q^{\theta}_{i}(\tilde{x}_{i}^{k},\tilde{x}_{\text{p}(i)})]$, which are unbiased (williams92_reinforce). Statistical significance of differences in performance is computed by comparing the area under the curve (AUC) of $f(x)$-sample mean over iterations curves, using paired two-sided t-tests over random seeds.
4.1 Comparison on synthetic functions
@@
Figure 2: Comparison of a naive EDA to DADO on synthetic problems. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$, b, $L=50$, and c, $L=200$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.
Here we used junction trees with meta-variable nodes containing only one original variable, i.e., an exact tree decomposition. We did so because this allowed us to better control the difficulty of the synthetic functions in that we could specify each edge component function as a table of size $D^{2}$. That is, an edge component function in which each node has only a single variable and thus can take on $D$ states each, is fully specified by a value for every possible pair of node choices, i.e., a table. For larger meta-variables, we would have required a more compact representation than a table (which scales exponentially with the number of variables in each meta-variable) such as a neural network-based function. However, randomly initialized neural networks tend to be overly smooth and unrealistically easy to optimize, and neural network outputs cannot generally be set in a controlled manner across the entire input domain due to its combinatorial size, making it nontrivial to obtain realistically difficult functions. We simulated one $f(x)$ for each of three sequence lengths, $L=\{25,50,200\}$, by first randomly sampling a junction tree topology, and then randomly specifying the component functions. Each node function, a $D$-vector, was sampled $f_{i}\in\mathbb{R}^{D}\sim\mathcal{N}(0,0.01)$, where $\mathcal{N}$ denotes a Gaussian distribution. Each edge function, a $D\times D$ matrix, was sampled, $f_{i,j}\in\mathbb{R}^{D\times D}\sim\mathcal{N}(0,0.0025)$. To ensure a reasonable degree of non-smoothness in $f(x)$, we further explicitly added what in biology is known as reciprocal sign epistasis (starr16_epistasis_evolution; li24_mlde_epistasis). Specifically, for each edge function we, twice, randomly assigned one of the 20 alphabet letters to each node, i) $x_{i}:=A,x_{j}:=B$ and ii) $x_{i}:=C,x_{j}:=D$. Next we sampled an effect size, $\lambda\sim\mathcal{N}(0,4)$. Finally, we set $f_{i,j}(x_{i}=A,x_{j}=B)=0$ and $f_{i,j}(x_{i}=C,x_{j}=B)=\lambda$, and also half of the time, $f_{i,j}(x_{i}=A,x_{j}=D)=0$ and $f_{i,j}(x_{i}=A,x_{j}=D)=\lambda$. Initial search distribution weights were drawn from $\mathcal{N}(0,0.0004)$ and the biases set to 0.
Both methods draw $K=100$ samples from their search distribution at each iteration. For all three sequence lengths, DADO outperforms the naive EDA, both visually, and in terms of a statistical test (Fig. 2). The final iteration is more concentrated in regions of high $f(x)$, and DADO also reaches these regions more quickly than the naive EDA, which we quantify with a statistical test on the area under the curve of the sample means. The difference is more pronounced for larger $L$, corresponding to larger design spaces. When we increase the number of samples drawn at each iteration to $K=1000$, both methods perform even better (Sec. A.6.1; Fig. A2).
4.2 Comparison on protein property predictors
@@
Figure 3: Comparison of a naive EDA to DADO on protein property predictive models. For each of four proteins of varying length, a, Amyloid, b, AAV, c, GB1, and d, TDP-43, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used a naive EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from two-sided paired t-tests that AUC of the per-iteration mean is different between methods, using the 20 mean curves.

### v2

@@ 2.1 Formal Exposition of DADO
The goal of DADO is to obtain a generative model, $p_{\theta}(x)$, that maximizes the EDA objective, $\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]$, while leveraging decomposability in $f(x)$ for optimization efficiency. DADO requires as input a decomposed version of $f(x)$, which can be described by an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. As noted in the introduction and shown in our experiments on proteins, identifying useful decomposability is feasible in practice. Given the junction tree topology, we write $f(x)=\sum_{i\in\mathcal{N}}f_{i}(\tilde{x}_{i})+\sum_{(i,j)\in\mathcal{E}}f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$, where we refer to $f_{i}(\tilde{x}_{i})$ and $f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})$ respectively as node and edge component functions (Fig. 1a); these are intimately related to the “epistatic landscape” of a protein property function (Sec. A.1.1). When the component functions are not known a priori, they can be parameterized and fit to labeled data.
We will begin by recalling how to do decomposition-aware exact (non-distributional) optimization, that is, to solve $\arg\max_{x}f(x)$. This problem is efficiently solved with a classical message-passing algorithm, which coordinates local optimizations across parts of the junction tree to obtain a single global optimum. Its efficiency comes from breaking optimization over all variables jointly into separate optimizations for each (smaller) meta-variable. Having loaded the reader with this intuition, we then adapt these ideas to distributional optimization, yielding DADO.
@@ 2.1.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models (Pearl, 1988; Shah, 2014), it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ (Vlassis et al., 2004). Given a function defined on an undirected junction tree, $\mathcal{T}$, message-passing makes use of its topology to find a global maximum. In particular, one roots the undirected junction tree to obtain a directed tree, inducing a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
To obtain the rooted tree, $\mathcal{T^{\prime}}:=(\mathcal{N^{\prime}},\mathcal{E^{\prime}})$, from the junction tree, one keeps the same nodes, $\mathcal{N^{\prime}}=\mathcal{N}$, chooses a root node, $r$, and directs all edges in $\mathcal{E}$ outward from $r$, yielding directed edges, $\mathcal{E^{\prime}}$. Although rooting at any node will suffice, we choose $r$ such that $\mathcal{T^{\prime}}$ has the shortest height possible. Message-passing finds a global optimum in two passes through the tree: one round of passing messages up from leaves to root and one round passing messages back down to the leaves.
Given $\mathcal{T^{\prime}}$, classical message-passing first uses dynamic programming to accumulate information from the leaves up to the root. Similarly to any dynamic programming procedure, we accumulate solutions to increasingly larger intermediate sub-problems. In this case, each sub-problem is to find the value of $f(x)$ evaluated on only a subset of meta-variables, rather than on the full set of variables in $x$. Each sub-problem is tractable owing to the decomposition of the objective function into component functions for each node and edge, and by respecting the partial order of sub-problems induced by $\mathcal{T^{\prime}}$. Specifically, one computes a value function, $V^{\text{max}}_{i}(\tilde{x}_{\text{p}(i)})$, for each node $i\in\mathcal{N^{\prime}}\setminus\{r\}$, which tells us for each setting of its parent, $\tilde{x}_{\text{p}(i)}$, the value of the intermediate objective function defined by the edge component function, $f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})$, plus all component functions over the sub-tree rooted at $i$, given that all nodes maximize their respective intermediate objectives. Computing value functions constitutes the first pass through the tree, from leaves to root:
@@
$Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}):=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+$ $Q^{\text{max}}_{i}(\tilde{x}_{i})\text{\;and\;}Q^{\text{max}}_{i}(\tilde{x}_{i}):=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\text{max}}_{c}(\tilde{x}_{i}).$
@@ 2.1.2 From classical message passing to distributional optimization
In the same way that an EDA transforms $\arg\max_{x}f(x)$ into a distributional optimization problem, we can rewrite the equivalent message-passing objective in Eq. 3 as DO. Because the original optimization problems over $x$ are equivalent, their DO formulations are equivalent too,
$\arg\max_{\theta}\mathbb{E}_{p_{\theta}(x)}[f(x)]=\arg\max_{\theta}\Big(\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\Big),$ (4)
@@
Classical message-passing (Eq. 2) independently maximizes each $Q$-function conditional on the choice of $\tilde{x}_{\text{p}(i)}$, instead of explicitly maximizing all variables in $x$ jointly (Eq. 3). This is possible because each $Q_{i}$ captures all relevant global information needed to choose $\tilde{x}_{i}$. It stands to reason then that DO can do something similar. Specifically, instead of training a single joint search distribution, we train smaller search distributions over each $\tilde{x}_{i}$, conditional on $\tilde{x}_{\text{p}(i)}$, to separately maximize each $Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})$. That is, we factor the search distribution according to $\mathcal{T^{\prime}}$, resulting in a DAG, $p_{\theta}(x)\coloneqq p_{\theta}(\tilde{x}_{r})\prod_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})$, for root node $r$, non-root nodes $i$, and parents $\text{p}(i)$, connected by directed edges $\mathcal{E^{\prime}}$, and with parameters, $\theta$. This factorized search distribution can be plugged into Eq. 4 for an equivalent optimization problem. We refer to each element of this product as one of the factors of the search distribution. In our implementation, each factor has completely separate parameters, though we write a shared $\theta$ for conciseness. Notice that each factor of the search distribution interacts with each other factor through the directed edges, $\mathcal{E^{\prime}}$. That is, the distribution of $\tilde{x}_{i}$ depends on its parent’s factor, and through it, all of its parent’s ancestors: $p_{\theta}(\tilde{x}_{i})=p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})$. In turn, each of node $i$’s children’s factors depends on $p_{\theta}(\tilde{x}_{i})$. Due to this coupling, we cannot optimize each factor fully independently. But we can still update each factor separately from the others in a globally-consistent manner via message-passing. In particular, each factor will only be responsible for directly optimizing its own meta-variable, but will need to coordinate with its neighboring factors by getting information from them about how they are optimizing their meta-variables in a manner analogous to classical message-passing. Our current messages, $Q_{i}^{\text{max}}$, convey the value of each intermediate objective when all meta-variables are chosen via maximization. For DO, we’ll require messages that communicate values when meta-variables are chosen according to the factorized search distribution, $p_{\theta}(x)$.
@@
$V^{\theta}_{i}(\tilde{x}_{\text{p}(i)})$ $:=\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})],\text{\;\;where}$
$Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)}):=f_{\text{p}(i),i}(\tilde{x}_{\text{p}(i)},\tilde{x}_{i})+$ $Q^{\theta}_{i}(\tilde{x}_{i})\text{\;and\;}Q^{\theta}_{i}(\tilde{x}_{i}):=f_{i}(\tilde{x}_{i})+\sum\nolimits_{c\in\text{children}(i)}V^{\theta}_{c}(\tilde{x}_{i}).$
@@
$\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\text{max}}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})]\geq\\ \mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(x)}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$ (5)
@@
$\arg\max_{\theta}\mathbb{E}_{p_{\theta}(\tilde{x}_{r})}[Q^{\theta}_{r}(\tilde{x}_{r})]+\sum\nolimits_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}\mathbb{E}_{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})p_{\theta}(\tilde{x}_{\text{p}(i)})}[Q^{\theta}_{i}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})].$
@@
$\theta^{n+1}=\arg\max_{\theta}\!\sum_{x\sim p_{\theta^{n}}(x)}\Big(Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta}(\tilde{x}_{r})+\!\!\sum_{(\text{p}(i),i)\in\mathcal{E^{\prime}}}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})\Big),$
which is equivalent to separate updates because the factors don’t share parameters, as desired:
$\theta_{r}^{n+1}$ $=\arg\max_{\theta_{r}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{r}^{\theta^{n}}(\tilde{x}_{r})\log p_{\theta_{r}}(\tilde{x}_{r})\text{, \;\; and}$
$\theta_{i}^{n+1}$ $=\arg\max_{\theta_{i}}\sum\nolimits_{x\sim p_{\theta^{n}}(x)}Q_{i}^{\theta^{n}}(\tilde{x}_{i},\tilde{x}_{\text{p}(i)})\log p_{\theta_{i}}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)}),\;\;\forall\;i\in\mathcal{N}\setminus\{r\}.$
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. We emphasize that DADO’s update is more statistically efficient than a naive EDA’s because DADO gets to use all $K$ samples for weighted MLE on each lower-dimensional factor distribution (i.e., ratio of number of samples to number of dimensions is larger). Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics (Brookes et al., 2020), included in Alg. 1. Choosing $W$ to be the identity recovers our derivation. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 5) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.5).
@@
1:Given: $\text{junction tree}\;\mathcal{T}=(\mathcal{N},\mathcal{E}),\;\text{component functions}\;\{f_{i,j}(\tilde{x}_{i},\tilde{x}_{j})\}_{(i,j)\in\mathcal{E}}\;\text{and}\;\{f_{i}(\tilde{x}_{i})\}_{i\in\mathcal{N}}$
@@
[…]
14:  for all $i\in\mathcal{N}$, in parallel do
@@ 3 Related Work
The closest work to ours in the EDA community is that of the Factorized Distribution Algorithm (FDA), wherein a decomposition is leveraged to factorize the conditional probability table (CPT) that parameterizes the search distribution (Mühlenbein and Mahnig, 1999). However, FDA cannot leverage message passing to give different weights to different variables so as to more efficiently update the search model. In related work, Pelikan (2005) replaced the CPTs with Bayesian networks, but still had no means to coordinate among variables through message passing or the like.
In a complementary line of work on policy optimization in reinforcement learning (analogous to learning the search distribution in EDAs), coordination between variables is enabled by message passing on a factorized search distribution/policy, but this line of work is only suitable for graph topologies that are Markov chains—that is, chain graphs. These graphs arise from reward functions, $f(x)$, that decompose in time according to a first-order Markov assumption (Peters and Schaal (2007), Peng et al. (2019), Nair et al. (2020)). Such approaches cannot be used on arbitrary junction trees.
Adjacent to our problem of interest, because they focus solely on real-valued design spaces, Grudzien et al. (2024) introduces a new way to discover a functional decomposition from data, and then demonstrate that doing so helps improve out-of-distribution generalization. Separately, the Bayesian optimization (BO) community has developed methods for dynamically inferring a function decomposition from data during active learning. Although they make use of message passing for coordination, they do not employ distributional optimization, nor can these methods be readily generalized to do so (Kandasamy et al. (2015), Han et al. (2021), Hoang et al. (2018), Rolland et al. (2018), Bardou et al. (2024)). This community has shown that using approximate or even random decompositions can be helpful (Ziomek and Ammar, 2023). DADO could, in principle, be used within the BO inner loop, although such an investigation is beyond the scope of the present work.
@@ 4 Experimental Results
We perform two sets of experiments with increasing resemblance to real-world scientific design. In each setting we compare a standard EDA, that is unaware of the function decomposition, to DADO, which is aware of it. In the first setting, we create synthetic functions $f(x)$, while in the second setting, we focus on functions derived from real protein data. In all experiments we designed sequences of fixed length $L$, where each position is one of $D=20$ amino acids. We used $N=100$ EDA training iterations and $G=1$ gradient steps for each weighted maximum likelihood update. We use MLP-based search distributions (details in Sec. A.8.1) for DADO and all baselines. We also compare to FDA (Mühlenbein and Mahnig, 1999) and PPO (Schulman et al., 2017) (details in Sec. A.7). For the shaping function we choose ${W}(s)=\exp\frac{s}{\beta}$ and tune the temperature, $\beta$. To pick $\beta$ and the learning rate, $\eta$, we performed a hyperparameter sweep separately for each method and each $f(x)$, taking $(\eta^{*},\beta^{*})$ with the largest sample mean at the last iteration. For synthetic experiments, we swept over 100 combinations between $\eta\in[10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.1,\;8]$. For protein experiments, which take longer, we swept only 54 pairs between $\eta\in[5\times 10^{-5},\;5\times 10{-2}]$ and $\beta\in[0.01,\;5]$. If either method’s $\eta^{*}$ or $\beta^{*}$ was on the boundary of the sweep, we broadened it. $(\eta^{*},\beta^{*})$ were then used for all replicate runs over random seeds dictating the initial search distribution and sampling. Although not essential to our method as described in Alg. 1, in our implementation we reduce the variance of the search distribution update by using mean-shifted $Q$-functions, $Q_{i}(\tilde{x}_{i},\tilde{x}_{p})-\mathbb{E}_{\{\tilde{x}_{i}^{k}\}}[Q_{i}(\tilde{x}_{i}^{k},\tilde{x}_{p})]$, which are unbiased (Williams, 1992). Statistical significance of differences in performance is computed by comparing the sample mean $f(x)$ at the final iteration, using paired two-sided t-tests over random seeds.
@@
Figure 2: Comparison of a naive EDA to DADO on synthetic problems. We created three random functions, $f(x)$, each with a randomly chosen junction tree decomposition with maximum node size of one, and randomly chosen parameters. Each experiment used alphabet size $D=20$ and sequence length, a, $L=25$, b, $L=50$, and c, $L=200$. Each of the two methods drew $K=100$ samples per iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 100 samples evaluated on $f(x)$, averaged across results from 20 random seeds. P-values are from a two-sided paired t-test that the mean at the final iteration is different between DADO and the most competitive baseline, over the 20 seeds.
Comparison on Fully Synthetic Functions
Here we used junction trees with meta-variable nodes containing only one original variable, i.e., an exact tree decomposition. We did so because this allowed us to better control the difficulty of the synthetic functions in that we could specify each component function as a dense lookup table of size $D^{2}$ and thereby ensure that there would be non-smoothness all throughout the design space. In contrast, larger nodes would have increased lookup table sizes exponentially with the number of variables in each meta-variable, requiring a more compact representation. Sparse lookup tables are one such option but generally result in needle-in-a-haystack-like functions. Neural networks are another, but random initialization tends to yield overly smooth functions that are unrealistically easy to optimize. Nor is it obvious how to produce neural landscapes that are realistically difficult across the entire design space via training. We simulated one $f(x)$ for each of three sequence lengths, $L=\{25,50,200\}$, by randomly sampling a junction tree and component functions (details in Sec. A.8.2). Search model weights were initialized from $\mathcal{N}(0,0.0004)$ and biases were set to 0.
All methods drew $K=100$ samples from their search distribution at each iteration. For all three sequence lengths, DADO outperforms all three baselines (Fig. 2). When $L=25$, the baselines catch up to DADO by the end of 100 iterations, but for larger $L$, corresponding to larger design spaces, the baseline methods converge to substantially lower values of $f(x)$ than DADO. Results from experiments in which we increase $D$ further support DADO’s superiority (Fig. A6).
@@
Figure 3: Comparison of a naive EDA to DADO on protein property predictive models. For each of four proteins of varying length, a, AAV, b, Amyloid, c, Gcn4, and d, TDP-43, we fit a neural network property function, $f(x)$, adhering to a junction tree decomposition derived from the protein’s 3D structure, and then used a naive EDA and DADO to optimize them. Each approach drew $K=1000$ samples per EDA iteration. For each iteration, we show the mean (solid line) and 95% confidence interval (shaded envelope) of the 1000 samples evaluated on $\log(c-f(x))$, averaged across results from 20 random seeds. We plot this quantity to make clear the differences between methods when $f(x)$ is large; $c$ is the largest $f(x)$ on a given plot, plus a small constant for numerical stability. P-values are from a two-sided paired t-test that the mean at the final iteration is different between DADO and the most competitive baseline, over the 20 seeds.
Comparison on Protein Property Predictors